# Steering Dialogue Dynamics for Robustness against Multi-turn Jailbreaking Attacks

### v1

@@ Appendix C Limitations and Future Work
Despite the effectiveness of our proposed NBF-based safety steering framework in mitigating multi-turn jailbreaking attacks, several limitations remain. First, our approach relies on learned state-space representations, which may not fully capture the complexities of language dynamics across diverse LLM architectures. While we demonstrate strong generalization to unseen models, performance may degrade when applied to models with significantly different architectures or training data distributions. Second, our method depends on high-quality labeled safety data, which can be costly and time-consuming to obtain. The effectiveness of the safety predictor and NBF is contingent on the quality and diversity of the training data, and biases in the dataset could affect performance. Third, while our method significantly improves safety, it introduces a trade-off with helpfulness. Higher steering thresholds ($\eta$) improve robustness against attacks but can lead to overly restrictive filtering, reducing the model’s ability to provide useful responses. Future research can work on adaptive steering mechanisms that dynamically adjust $\eta$ based on conversational context and user intent. Finally, our approach assumes that attack queries follow known multi-turn jailbreaking strategies. While we show some resilience to adaptive attacks (e.g., synonymic reformulations), stronger adversarial strategies could be developed to circumvent our filtering mechanism. Future research could explore adversarial training and online adaptation to enhance robustness against evolving attack patterns.

### v2

@@ Appendix C Limitations and Future Work
Despite the effectiveness of our proposed NBF-based safety steering framework in mitigating multi-turn jailbreaking attacks, several limitations remain. First, our approach relies on learned state-space representations, which may not fully capture the complexities of language dynamics across diverse LLM architectures. While we demonstrate strong generalization to unseen models, performance may degrade when applied to models with significantly different architectures or training data distributions. Second, our method depends on high-quality labeled safety data, which can be costly and time-consuming to obtain. The effectiveness of the safety predictor and NBF is contingent on the quality and diversity of the training data, and biases in the dataset could affect performance. Third, while our method significantly improves safety, it introduces a trade-off with helpfulness. Higher steering thresholds ($\eta$) improve robustness against attacks but can lead to overly restrictive filtering, reducing the model’s ability to provide useful responses. Future research can work on adaptive steering mechanisms that dynamically adjust $\eta$ based on conversational context and user intent. Finally, our approach assumes that attack queries follow known multi-turn jailbreaking strategies. While we show some resilience to adaptive attacks (e.g., synonymic reformulations), stronger adversarial strategies could be developed to circumvent our filtering mechanism. Future research could explore adversarial training and online adaptation to enhance robustness against evolving attack patterns. Regarding the broader impacts, as a post-training method, the proposed LLM steering can be extended to agentic applications beyond safety, where the LLM agent should focus on specific topics and avoid other topics in the multi-turn conversation settings.